OpenStudy (anonymous):

Find the circumference of the circle pictured. Use C = 3.14(2)(r). A. 6.28x + 3.14 B. 3.14^2 + 6.28 C. 12.56x + 6.28

OpenStudy (anonymous):

need picture

OpenStudy (anonymous):

do you have a picture or can you tell the diameter/radius?

OpenStudy (anonymous):

oh sorry..lol i forgot..hold on i'll post it

OpenStudy (anonymous):

r = 2x +1 c =3.14 (2x + 1)(2) c = 6.28 (2x +1) c = 12. 5 6x + 6.28

OpenStudy (anonymous):

they gave you the radius - so plug into your formula for C and solve

OpenStudy (anonymous):

oh okay thnk u J..ur a big hlp

OpenStudy (anonymous):

you are welcome